OpenStudy (avanti):

find the 4th roots of the complex number z1=1+sqrt3i part1: write z1 in polar form part2: find the mudulus of the root of z1 part3: find the four angles that define the 4th roots of the number z1 part4: what are the fourth roots of z1=sqrt3+i

OpenStudy (anonymous):

....

OpenStudy (psymon):

Well, polar form means you need: \[r(\cos \theta +isin \theta)\] r is the modulus and theta is the argument. There are several conversions to go back and forth between rectangular and polar, but we need these two: \[r = \sqrt{a ^{2}+b ^{2}}\] \[\tan \theta =\frac{ b }{ a }\] Now since these come in the form of "a+bi", whatever is not paired with i will be your "a" and what is with i is your "b". So it looks like you have: \[a = 1 + \sqrt{3}\] \[b = 1 \] So using that as your a and b and using the conversion above, you think you could get r?

OpenStudy (avanti):

this may seem like really basic math but how do i square 1 + sqrt(3) ?

OpenStudy (psymon):

\[\sqrt{(1)^{2}+\sqrt{3}^{2}} = \sqrt{1+3}\]

OpenStudy (psymon):

Hmm....wait, nvm.

OpenStudy (psymon):

I did that without thinking, haha.

OpenStudy (psymon):

Well, we can either turn it into a decimal so we can combine it or we have to do this: \[r=\sqrt{(1+\sqrt{3})^{2}+1}\] And then its up to you what yoiu want to do with that. But that is sqrt(a^2 + b^2). Do you want to simplify it, make it a decimal, or what? It is kind of weird since you have a problem like this, though xD

OpenStudy (avanti):

i just made it a decimal. I used my calculator and i got 7.464101615...

OpenStudy (avanti):

for (1+sqrt(3)) ^2

OpenStudy (psymon):

Ah. Alright, I'll do the whole a^2 + b^2 calculator work.

OpenStudy (psymon):

So 2.91

OpenStudy (avanti):

Okay so I'm getting r to equal=2.9 right?

OpenStudy (avanti):

oh yeah haha we both got it!

OpenStudy (psymon):

Are you sure theres a plus between sqrt 3 and i?

OpenStudy (avanti):

oh no there is not! haha sorry my fault

OpenStudy (psymon):

Okay, no wonder x_x Makes this much easier. Okay, so then our r is this: \[\sqrt{(1)^{2}+(\sqrt{3})^{2}}\]

OpenStudy (psymon):

Which I actually posted above, lol.

OpenStudy (avanti):

Yes! so r=2

OpenStudy (psymon):

Yeah xD Much more sense. R = 2 and r is the modulus. So now we need theta. We do that using this conversion: \[\tan \theta = \frac{ b }{ a }\] Think you might know how to get that?

OpenStudy (avanti):

yes theta = sqrt(3)?

OpenStudy (psymon):

Tan(theta) = sqrt(3). You only have tangent, you need the actual theta xD

OpenStudy (avanti):

oh that means theta=60 degress!

OpenStudy (psymon):

Sounds good to me, lol. I guess you need it in degrees. So what if it were 1 - sqrt(3)i. Then what would theta be? (just to make sure we get the idea)

OpenStudy (avanti):

wouldn't it be the same thing since -sqrt3 * -sqrt3 is still 3?

OpenStudy (psymon):

Thats if we find r, though, not theta. Theta would definitely be different.

OpenStudy (avanti):

hmm I'm not exactly sure.. 300 degrees?

OpenStudy (psymon):

Yep xD The reason I ask is because you almost always havetwo options for theta. tan(theta) = sqrt(3) has two answers, but you need to know which is the correct one. The way you know is to look at your original rctangular problem. With us, we need to go right 1 and then up sqrt(3) units. In the example I gave you, itd be right 1 but down sqrt(3) units. So I just wanted to bring that to your attention :P

OpenStudy (avanti):

okay that's good to know! Do you know how to do part 3 or 4 of this question?

OpenStudy (psymon):

Well, 3 and 4 are almost the same problem. But okay, so this is what we should have so far for our polar form: \[2(\cos60 + isin60)\]Now in order to find roots, we have to use this formula: \[\sqrt[n]{r}(\cos(\frac{ \theta }{ n }+\frac{ 2k \pi }{ n })+isin(\frac{ \theta }{ n }+\frac{ 2k \pi }{ n }))\] Where n is theroot we need. So basically we take the nth root of r and then divide each angle by whatever root what we want. After we do that, we have our first root. To get the other roots, you use the + 2kpi/n part. So if we need a 5th root, think you can find the first of the 5 5th roots?

OpenStudy (avanti):

Yikes this math is way beyond me... I don't really know :(

OpenStudy (psymon):

Well, we have \[2(\cos60 + isin60)\] So just take the 5th root of r and divide the angles by 5.

OpenStudy (avanti):

so since r is 2, we take the 5th root of 2? which is apprx 1.15 ?

OpenStudy (psymon):

Well, you usually don't put it into decimals unless it asks you to. Decimals we usually try to avoid if we can xD I would just leave it as: \[\sqrt[5]{2}\] or something like that xD

OpenStudy (avanti):

okay :) are you supposed to divide 5th root of 2 by 4 to get part 3?

OpenStudy (psymon):

Oops, we needed 4th roots, lol. My bad x_x

OpenStudy (psymon):

\[\sqrt[4]{2}\]

OpenStudy (psymon):

So thats the radius of our 4th roots. Now divide the angles by 4.

OpenStudy (avanti):

which angles? :o

OpenStudy (psymon):

Did you never get or see the polar form we came up with? :P

OpenStudy (avanti):

oops sorry is it cos60 and sin60?

OpenStudy (psymon):

Yeah xD So divide those by 4.

OpenStudy (avanti):

so 1/8 and sqrt(3)/8

OpenStudy (psymon):

If you turn them into that, then yes, lol. Can you put all of that in standard form before we continue?

OpenStudy (avanti):

1/8 + sqrt(3)/8 i

OpenStudy (psymon):

No r? xD

OpenStudy (avanti):

oops im getting really mixed up.. i dont know where the r goes.

OpenStudy (psymon):

You just dont remember the standard form is all :P \[r(\cos \theta + isin \theta)\]

OpenStudy (avanti):

okay 2(cos(60) + i sin(60) = 2(1/8 + sqrt(3)/8 i)

OpenStudy (psymon):

Sorry to keep making corrections here and there, but we have to keep it in sin and cos, we cant say cos60 = 1/2 or anything, we'll getthe wrong answers. You just gotta get into the habit of not making everything decimals or simplifying things unless we need to xD Sorry @_@

OpenStudy (avanti):

oh its okay! so what are the 4 angles?

OpenStudy (psymon):

Well, whats 60/4? :P

OpenStudy (avanti):

15

OpenStudy (psymon):

Right. So this is the first answer. \[\sqrt[4]{2}(\cos15+isin15)\] Now you gotta now how we get here, though. We've spenta lot of time, somake sure you know what we did to get here o.o

OpenStudy (avanti):

yeah i think i understand it now! would you just add another 15 degrees to get the 2nd angle?

OpenStudy (psymon):

Nope. Remember thewhole 2kpi/n thing?

OpenStudy (avanti):

oh right so add 360 degrees?

OpenStudy (psymon):

uh, no o.o you add: \[\frac{ 2k \pi }{ n }\] where n is the root you want. So sincen is 4, we would add: \[\frac{ k \pi }{ 2 }\] where k is 0, 1, 2, 3. k = 0 was our first 4th root that we just typed. Adding that when k = 1 is the 2nd, adding that when k = 2 is the 3rd and k = 3 is the 4th. Although Im sure ill have to explain and show more than that, lol.

OpenStudy (avanti):

hmm okay thank you for all your help, i really really appreciate it! :D

OpenStudy (psymon):

Lol, you think you got it?

OpenStudy (avanti):

no but its really late and i dint think im gonna get it but thanks again for all your efforts! i understand it better than i did!

OpenStudy (avanti):

*don't

OpenStudy (psymon):

Well we'll get it later then, lol.

OpenStudy (avanti):

yeah thank you for trying! :)

OpenStudy (psymon):

Night, lol.